OpenStudy (anonymous):

Last one -5/8t=4/5 Solve ?

OpenStudy (anonymous):

-5/8t = 4/5 -5/t = 4/5 * 8 -5/t = 32/5 1/t = 32/5 * -1/5 1/t = -32/25 t = -25/32

OpenStudy (anonymous):

Thank you so so much!

OpenStudy (anonymous):

hold on i messed up lol

OpenStudy (anonymous):

nvm i was right hahah